OpenStudy (anonymous):

Find the distance between the points (-1, 3) and (5, 0). (Hint: Remember to simplify the square root if you can. For example: sqrt 28 = 2 sqrt 7.) please show steps....

OpenStudy (anonymous):

3 sqrt 5

OpenStudy (anonymous):

Apply the distance formula again: \[D=\sqrt{9+36}=\sqrt{45}=3\sqrt{5}\]

OpenStudy (anonymous):

sqrt((5-(-1))^2 + (0-3)^2) sqrt(6^2 + 3^2) sqrt(36+9) sqrt(45) 3√5

OpenStudy (anonymous):

d^2 = (x1-x2)^2 + (y1-y1)^2 x1,y1 = 5,0 x2,y2 = -1,3 d^2 = (5-(-1))^2 + (0-3)^2 d^2 = 6^2 + -3^2 d^2 = 36 + 9 d = 45^0.5 d = (9*5)^0.5 d = (9^0.5)*(5^0.5) d = 3 * (5^0.5)

OpenStudy (amistre64):

the steps dont change from distance problem to distance problem ... that you can be assured of

OpenStudy (amistre64):

subtract the points and pythag them